Jessica is 4 years younger than Nadia.
8 years ago, Nadia was 3 times as old as Jessica.

Answer by VFBundy(438)

J = Jessica's age
N = Nadia's age

Jessica is 4 years younger than Nadia:
J = N - 4

Put all the variables on the left side of the equation:
J - N = -4

8 years ago, Nadia was 3 times as old as Jessica:
N - 8 = 3(J - 8)

Simplify above:
N - 8 = 3J - 24

Put all the variables on the left side of the equation:
-3J + N = -16

List both equations:
J - N = -4
-3J + N = -16

Add both equations together:
-2J = -20

Solve for J:
J = 10

Substitute J = 10 into any of the equations to find N:
J = N - 4

10 = N - 4

N = 14

J = Jessica's age = 10
N = Nadia's age = 14